# Metis-SPECS: Decoupling Learning via Self-distilled Preference-based Cold Start for VLMs

### v2

@@ Appendix F Experiments Implementation Details
All experiments were conducted on a cluster of 32 NVIDIA H800 (80G) GPUs. We utilize the open-source Multimodal Large Language Model, Qwen2.5-VL-7B (Qwen2.5-VL), as our base model.
For the GRPO training in Stage 1 and Stage 3, we employ the MM-EUREKA  [^7] https://github.com/ModalMinds/MM-EUREKA framework. The rollout and training batch sizes are both set to $128$, with $8$ rollouts generated per sample. The learning rate is configured to $1\times 10^{-6}$. Following the DAPO (yu2025dapo) methodology, we set the clipping thresholds to $0.2$ (lower) and $0.28$ (upper), and we do not apply a KL penalty. The maximum output length is restricted to $10,240$ tokens. This training phase spanned $400$ steps and was completed in approximately $12$ hours.

### v3

@@ Appendix F Experiments Implementation Details
All experiments were conducted on a cluster of 32 NVIDIA H800 (80G) GPUs. We utilize the open-source Multimodal Large Language Model, Qwen2.5-VL-7B (Bai et al., 2025), as our base model.
For the GRPO training in Stage 1 and Stage 3, we employ the MM-EUREKA  [^6] https://github.com/ModalMinds/MM-EUREKA framework. The rollout and training batch sizes are both set to $128$, with $8$ rollouts generated per sample. The learning rate is configured to $1\times 10^{-6}$. Following the DAPO (Yu et al., 2025) methodology, we set the clipping thresholds to $0.2$ (lower) and $0.28$ (upper), and we do not apply a KL penalty. The maximum output length is restricted to $10,240$ tokens. This training phase spanned $400$ steps and was completed in approximately $12$ hours.